OpenStudy (evansanchez):

Which two points are on the graph of y = -x + 3 (-1, -2), (1, 4) (1, 2), (0, -3) (0, 3), (4, -1) (4, -1), (1, 3)

OpenStudy (evansanchez):

is it C?

jimthompson5910 (jim_thompson5910):

Let's check Plug in (x,y) = (0,3). So we replace x with 0 and y with 3 y = -x + 3 3 = -0 + 3 3 = 3 ... works Since the equation is true, this means (0,3) is on the line Now plug in (x,y) = (4,-1). So we replace x with 4 and y with -1 y = -x + 3 -1 = -4 + 3 -1 = -1 ... works Since the equation is true, this means (4,-1) is on the line

jimthompson5910 (jim_thompson5910):

So you are correct